11.   There once was a lonely and rather sour buzzard named Clyde. It came as no surprise to any who knew him that Clyde had no offspring‑‑no female buzzard would come anywhere near a buzzard with his personality. Clyde, however, nursed a secret desire to pass on his genes, and one day he hit upon a plan. He had heard from his boss, Professor Johnson, that the St. Louis Zoo practices birth control among its captive birds by the simple expedient of keeping only female birds. In dark of night, he invaded the zoo and there wooed numerous female buzzards, none of whom knew the meanness of his nature. Soon buzzard eggs were hatching all over the zoo.

Now a new issue arose to give Clyde pain. In reading his boss's genetics notes, he made a dread discovery: genes recombine during meiosis. This meant that there was a chance that his wonderfully horrible combination of characteristics might be diluted out by other more "normal" alternatives in subsequent generations.

Clyde brooded on this for quite a while. But he finally decided that he need not worry, since he remembered his scrawny mother telling him on her knee that the two traits he most cared about, dumb mind and hard heart, were closely linked to one another; his mother was in fact homozygous for these traits, as well as for scrawniness of frame. Because Clyde shows all three traits, even though his father was normal and did not, all three traits are dominant.

Clyde asked his boss to look into this matter for him by examining the baby buzzards at the zoo. Dr. Johnson located 1000 baby buzzards apparently sired by Clyde in his nightly visits and checked out the dumbness of their minds, the hardness of their hearts, and the scrawniness of their frames. Here is what he found:

NUMBER   MIND       FRAME       HEART

235   Normal       Normal       Normal

230   Dumb       Normal       Hard

226   Dumb       Scrawny   Hard

221   Normal       Scrawny   Normal

24   Dumb       Normal       Normal

23   Normal       Normal       Hard

21   Normal       Scrawny   Hard

20   Dumb       Scrawny   Normal


Dr. Johnson went home and told Clyde he had nothing to worry about, that hardness of heart and dumbness of mind are very closely linked, relative to other Clyde traits such as scrawniness of frame. Was he right? Explain, backing up your argument with an appropriate chromosome map.

bi\MAJORS\211\ExtraCredit\GeneMaps   adapted from G. Weber (rev7/25/94)

PLEASE HELP I HAVE NO IDEA HOW TO DO ANY OF THIS?? Can I get this in depth so I can attempt this on exams :(
Answer #1

Answer:

NUMBER   MIND       FRAME       HEART

235   Normal       Normal       Normal (DSH)

230   Dumb       Normal       Hard (dSh)

226   Dumb       Scrawny   Hard (dsh)

221   Normal       Scrawny   Normal (DsH)

24   Dumb       Normal       Normal (dSH)

23   Normal       Normal       Hard (DSh)

21   Normal       Scrawny   Hard (Dsh)

20   Dumb       Scrawny   Normal (dsH)

Triple zygote = DSH/dsh

1).

If single crossover occurs between d&s

Normal combination: DS/ds

After crossover: Ds/dS

Ds progeny=221+21=242

dS progeny = 230+24=254

Total this progeny = 496

The recombination frequency between d&s = (number of recombinants/Total progeny) 100

RF = (496/1000)100 = 49.6%

2).

If single crossover occurs between s&h

Normal combination: SH/sh

After crossover: Sh/sH

Sh progeny= 230+23=253

sH progeny = 221+20 = 241

Total this progeny = 494

The recombination frequency between s&h = (number of recombinants/Total progeny) 100

RF = (494/1000)100 = 49.4%

3).

If single crossover occurs between d&h

Normal combination: DH/dh

After crossover: Dh/dH

Dh progeny=23+21=44

dH progeny = 20+24=44

Total this progeny = 88

The recombination frequency between d&h = (number of recombinants/Total progeny) 100

RF = (88/1000)100 = 8.8%

Recombination frequency (%) = Distance between the genes (mu)

d----8.8mu---h----------------------49.4mu-------------s

The order of genes = d-h-s

Based on this gene map we could conclude that hardness of heart and dumbness of mind are very closely linked, relative to other Clyde traits such as scrawniness of frame.